40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jul 25 10:48:13 2008, Koji, Update, General, Abs. Len. Meas. ~ Realignment / beating / PLL trial 
    Reply  Mon Jul 28 15:54:04 2008, Koji, Update, General, Abs. Len. Meas. ~ More on the beat / the PLL setup PLL_setup.pngbeat_at_8MHz.jpg
       Reply  Thu Jul 31 11:19:30 2008, Koji, Update, General, Abs. Len. Meas. ~ Resonance search trial 
          Reply  Sat Aug 2 12:53:43 2008, Koji, Update, General, Abs. Len. Meas. ~ New PLL at the PSL table yarm_dc.png
             Reply  Mon Aug 4 00:37:58 2008, Koji, Update, General, Abs. Len. Meas. ~ RF PD at the Y end / Manual frequency scan YEND_LAYOUT.png4th_FSR1.png4th_FSR2.png
                Reply  Mon Aug 4 05:23:57 2008, Koji, Update, General, Abs. Len. Meas. ~ Measurement for Y-arm completed 
                Reply  Mon Aug 4 21:48:24 2008, Koji, Update, General, Abs. Len. Meas. ~ Scan for TEM00/01/10 TEM00.pngTEM01.pngknife_edge.png
                   Reply  Mon Aug 4 22:31:10 2008, Koji, Update, General, Abs. Len. Meas. ~ Simple Test for TEM01/10 split TEM01_10.png
                   Reply  Tue Aug 5 00:05:57 2008, Koji, Update, General, Abs. Len. Meas. ~ IFR2023A calibration 
                      Reply  Wed Aug 6 11:10:34 2008, Koji, Update, General, Abs. Len. Meas. ~ analysis of the TEM00 scan TEM00fit.pngTEM00FSRfit.png
                         Reply  Wed Aug 6 11:43:52 2008, Koji, Update, General, Abs. Len. Meas. ~ analysis of the TEM01 scan TEM01fit.png
                            Reply  Thu Aug 14 21:52:51 2008, Koji, Summary, General, Abs. Len. Meas. ~ summary of my Summer mode_spacing_measurement_080816_v2.pdf
Message ID: 782     Entry time: Sat Aug 2 12:53:43 2008     In reply to: 776     Reply to this: 787
Author: Koji 
Type: Update 
Category: General 
Subject: Abs. Len. Meas. ~ New PLL at the PSL table 
Report of the work last night:
The new heterodyne interferometer on the PSL table was built.
The length of the Yarm cavity was measured with better precision.

-------------
Yarm is locked. The injection beam was aligned. The beat was there at around LT=48.9 [C_deg] of the NPRO.

The new PLL setup on the PSL table has been built. The two beams from the MC incident beam and the injection beam are
mode-matched with lenses. I measured the Rayleigh ranges of the beams by a sensor card and my eyes, and then placed
appropriate lenses so that they can have 5~6 [m] Rayleigh range. They looks a bit too thick but just ok for an inch
optics. The new PLL setup shows ~70% intensity modulation which is enormous. The servo is still SR560-based so far.

Now the PLL has no singular frequency within its range. I could sweep the 4th FSR of the cavity with 500Hz interval. I
was still observing at the transmitted DC.

At each freqency from 15.51MHz to 15.52MHz, a timeseries data of the Yarm transmitted was recorded at sampling of 32Hz for 10
seconds. The figure shows the averaged values of the transmitted DC with errors. An increase of the transmitted power by
3-4% was found. If we consider the resonance is at f_PLL = 15.515 +/- 0.0005 [MHz], this indicates the
arm cavity length of 38.6455 +/- 0.0012 [m].
Y-arm length
e-log    length [m]
-----------------------------
556      38.70   +/- 0.08   Cavity swinging measurement
556      38.67   +/- 0.03   tape & photo
776      38.640  +/- 0.007  Beam injection, poor PLL, Transmitted DC
this     38.6455 +/- 0.0012 Beam injection, independent PLL, Transmitted DC
-----------------------------

NEXT STEPS:
o RF detection at the transmitted
o Better PLL: PLL stability (in-loop / out-of-loop)
o Measurement for the 1st~3rd FSR
o Reproducibility of the measurement
o Higher order mode search
o Check the acuuracy and presicion of the Marconi
Attachment 1: yarm_dc.png  53 kB  | Hide | Hide all
yarm_dc.png
ELOG V3.1.3-